Use the base-change formula to find the logarithm to 4 decimal places.
log base 8 X 8.69

Answer by RAY100(1637)

log(8) of 8.69 = log 8.69 / log 8 = .939 / .903 = 1.040
.
or = ln 8.69 / ln 8 = 2.162/ 2.079 =1.040
.
remember,,, log to base a of b = log b / log a
.